OpenStudy (anonymous):

You have a jar of pennies and quarters. You want to choose 15 coins that are worth exactly $4.35 A. Write and solve a system of equations that models the situation B. is your solution reasonable in terms of the original problem? explain.

OpenStudy (anonymous):

please help me!!!!

OpenStudy (jdoe0001):

let us say p = "how many pennies there are" and q ="how many quarters there are" if you tip over the jar, and start counting them all, how many would you end up with?

OpenStudy (anonymous):

20

OpenStudy (jdoe0001):

heheh 20? well anyhow.... if you just hmmmm hold the mayo alright notice that 4.35 is what you'd want to choose

OpenStudy (anonymous):

yea...

OpenStudy (jdoe0001):

so, what is stated is that, you'd pick from the jar, 15 coins, quaters and pennies and those 15 coins will add up to $4.35 so, let's us see, $4.35 is really 435 pennies so if "p" is the quantity of penny coins and "q" is the quantity of quarter coins \(\Large \bf p + q = 15\) --------------------------------------------- there is 1 penny per penny, well that's obvious and there are 25 pennies per quarter how many pennies worth are there? well 1 * p how many pennies are there in the quarters? well 25 * q so 1*p + 25*q = 435 pennies or \(\Large \bf p+25q = 435\)